Demand function for a cola-type soft drink in general is q= 20 – 2P, where Q stand
for quantity and P stands for price.

a. Calculate point elasticities at prices of 5 and 9. Is the demand curve elastic or inelastic at these points?

b. Calculate arc elasticity at the interval between P = 5 and P = 6.
1
Expert's answer
2011-07-28T10:59:12-0400
Point elasticity (Ed).
Ed=δQ/δP * P/Q , where δQ/δP= -2
P1=5, Q1=20-2*5=10;
P2=9, Q2=2;
Point elasticity at the price P1: Ed1 = -2 * 5/10= -1. Unitarily elastic demand.
Point elasticity at the price P2: Ed2 = -2 * 9/2= -9. The demand curve is elastic.
Arc elasticity (Ed).
Ed=δQ/δP * (P1+P2)/(Q1+Q2 )

P1=5, Q1=10;
P2=6, Q2=8;
Ed1= -2/1 * 11/18= -11/9=-1.22
The demand curve is elastic.